Matt wants to buy a new riding lawn mower. The price is $2595. He finances the mower over 36 months after making a down payment of $395. The true annual interest rate is 12%. What is Matt’s monthly payments? (Principal plus interest)
Use the formula I = (2yc)/(m(n+1))
1
Expert's answer
2019-04-11T13:24:15-0400

I(Interest)= 2YC/(M(N+1)), where y = payments per year; c = total interest paid; m = amount financed; n = total number of payments.

0.12 = 2×12×C/((2595 - 395)×37) ,

C = 0.12×2200×37/(2×12) = $407.